Immunology report:

Final Bead Concentration

4.32 X 10^8

1.005 X 10^8

4.2 X 10^8

2.1 X 10^7

4.92 X 10^8

4.85 X 10^8

4.41 X 10^8

4.84 X 10^9

4.32 X 10^8

4.305 X 10^8

4.73 X 10^8

6.3 X 10^8

4.59 X 10^8

2.69 X 10^8

4.35 X 10^8

4.26 X 19^8

Q1) calculate mean bead concentration, standard deviation of the mean, and CV% bead concentration.

Q2) What can we say about the precision and accuracy of the bead count?
Answer #1

Q1)

Done using excel

Final Bead Composition (*10^8) 4.32 1.005 4.2 0.21 4.92 4.85 4.41 4.84 4.32 4.305 4.73 6.3 4.59 2.69 4.35 4.26 Mean 4.01875 *

so the mean is 4.01875*10^8 and the standard deviation is 1.509*10^8.

Coefficient of variation (CV) can be defined as standard deviation upon the mean.

ie, CV = standard deviation / mean

= (1.509*10^8)/(4.01875*10^8)

= 0.375

CV% = 0.375*100 = 37.5%

Q2)

Precision can be defined as the refinement of the value.Precision depends upon how many digits are there after the decimal point.

Accuracy can be defined as the closeness of the measured value to the real value.

here the values are less precise since only two digits are followed after the decimal point. We can say here the values are more accurate since the coefficient of variation (CV) is less. Only less variation is there from the mean. So the data is accurate.